OpenStudy (anonymous):

solve for x. 27^(2x)=9^(x-3)

OpenStudy (anonymous):

i am not playing with logs this early

OpenStudy (anonymous):

x = -3/2 Equalize the bases, and then solve for x.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No Problem.

OpenStudy (anonymous):

\[27^{2x}=9^{x-3}\]\[3^{6x} = 3^{2x-6}\] then just set \[6x = 2x -3\] No logs necessary!

OpenStudy (anonymous):

27^(2x)=9^(x-3) Create equivalent expressions in the equation that all have equal bases. 3^(3(2x))=3^(2(x-3)) Since the bases are the same, then two expressions are only equal if the exponents are also equal. 3(2x)=2(x-3) Multiply 2 by each term inside the parentheses. 3(2x)=(2x-6) Remove the parentheses around the expression 2x-6. 3(2x)=2x-6 Multiply 3 by each term inside the parentheses. 6x=2x-6 Since 2x contains the variable to solve for, move it to the left-hand side of the equation by subtracting 2x from both sides. 6x-2x=-6 Since 6x and -2x are like terms, add -2x to 6x to get 4x. 4x=-6 Divide each term in the equation by 4. (4x)/(4)=-(6)/(4) Simplify the left-hand side of the equation by canceling the common terms. x=-(6)/(4) Simplify the right-hand side of the equation by simplifying each term. x=-(3)/(2)

OpenStudy (anonymous):

Great explanation!

OpenStudy (anonymous):

thanks apupo 03